# PACEE: Supporting Children’s Personal Emotion Education through Parent-AI Collaboration

### v1

@@ Abstract.
Emotion education is a crucial lesson for children aged 3 to 6. However, existing technologies primarily focus on promoting emotion education from the child’s perspective, often neglecting the central role of parents in guiding early childhood emotion development. In this work, we conducted co-design sessions with five experienced kindergarten teachers and five parents to identify parental challenges and the roles that AI can play in family emotion education. Guided by these insights, we developed PACEE, an assistant for supporting parent-AI collaborative emotion education. PACEE enables parents to engage in emotional dialogues about common scenarios, with multiple forms of support provided by generative AI. It combines insights from parents and AI to model children’s emotional states and collaboratively delivers personalized, parent-mediated guidance. In a user study involving 16 families, we found that PACEE significantly enhances parent-child engagement, encourages more in-depth emotional communication, and improves the parental experience. Our findings advance emotion coaching theory in both family settings and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
@@
journalyear: 2025copyright: ccconference: ; ; $c$s: Human-centered computing Collaborative interaction
@@ 1. INTRODUCTION
Emotion education plays a critical role in the developmental years of children aged 3 to 6. According to developmental psychology, during this period, children begin to understand various emotions — such as anger, fear, and sadness — and learn to express and regulate these feelings (bartlett1982emotional; saarni1999development; wulandari2021development).
Recent studies have explored AI-based systems that help children recognize, express, and manage their emotions. However, these systems mainly target children aged 8 and above, as they rely on independent conversations with LLM-based agents — for example, expressing emotions through dialogue (seo2024chacha; tang2024emoeden; shen2025easel). Such designs are not well-suited for younger children aged 3 to 6, since at this stage, parents play a foundational role in shaping children’s emotional development by guiding their early emotional learning at home (fadlillah2022analysis; alegre2011parenting). Unlike older children, those aged 3 to 6 require parental support to engage with technology (toran2024parent; morawska2023managing) and rely on their parents for emotional security and co-regulation (corwin2012secure; brock2016interparental).
However, providing effective emotion education can be challenging for parents, as it requires knowledge of child psychology and strong communication skills — expertise most parents lack (sanders1999triple; saarni1999development; cole1994development). In practice, parents often encounter difficulties like misreading children’s ambiguous emotional signals (castro2015parents), or failing to support emotional regulation through dialogue (meng2020effects).
@@
We developed PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education for children aged 3 to 6. PACEE enables parents to engage in emotion education with their children through interactive adventure games. PACEE consists of three key modules, each addressing one of the aforementioned challenges: (1) Parent-Child Adventure Game. In each adventure game, parents and their children engage in five-step emotional dialogues based on Gottman’s Emotion Coaching Theory (gable1999heart). The AI system provides parents with real-time guidance, such as actionable suggestions and progress tracking, thus solving $C_{1}$. (2) Parental Feedback Report. This module provides parents with feedback after each adventure game. The feedback includes a comprehensive review of past parent-child conversations and actionable suggestions for future interactions, effectively addressing $C_{2}$. (3) Child Emotional Modeling. This module creates an emotional profile for each child by combining two views: long-term understanding provided by parents and AI’s analysis of parent-child dialogue data in the adventure game. The emotional profile is then used to offer tailored guidance to parents during the adventure game. After each adventure game, it also provides parents with a comparison between their own understanding of their child’s emotional states and AI observations, thereby improving parents’ understanding of their child and effectively addressing $C_{3}$.
To evaluate PACEE’s effectiveness in supporting family emotion education, we conducted a user study with 16 pairs of parent and child. Compared to a baseline (LLM-assisted) method, PACEE significantly fosters meaningful parent-child interactions and in-depth emotion education. Furthermore, with PACEE, children show greater engagement and richer emotional expression during conversations. Meanwhile, parents (1) improve their expertise and confidence in communication, (2) adopt more effective approaches in emotional education, and (3) gain a better understanding of their children’s emotional states.
@@
  • Through a formative study, we identified three key challenges in family emotion education, AI’s role in supporting this process, and principles for AI-assisted emotion education from experts’ perspectives.
  • We present PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education. PACEE collaboratively builds children’s emotional profiles with parents and provides personalized, parent-mediated guidance.
  • Through a user study with 16 families, we validated PACEE’s effectiveness in enhancing family emotion education and advanced emotion coaching theory in both family and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
@@ 2.1. Emotion Coaching Techniques and Theories
Emotion coaching is a supportive parenting technique that helps children understand and regulate emotions and has been widely studied in parenting and child development (fu2022social; havighurst2021tuning). A key framework widely applied in HCI (havighurst2021tuning) is the “Five Steps of Emotion Coaching” framework by John Gottman (gottman2005five), which includes (1) attending to children’s emotional signals, (2) viewing emotional moments as opportunities for connection and learning, (3) helping children identify and articulate emotions, (4) expressing empathy and validating emotions, and (5) providing guidance and boundaries while collaboratively exploring solutions. This model is often complemented by principles from positive psychology (baker2017positive; fan2018emostory), such as the PERMA model (seligman2010flourish). PERMA model outlines five elements of well-being (Positive Emotion, Engagement, Relationships, Meaning, and Achievement) and has been applied in HCI areas such as well-being-centered design (kovich2023application), persuasive technologies (nebrida2018m), and digital mental health (chisale2022perma).
Current HCI implementations of emotion coaching primarily target older children or specific clinical populations. Research has explored using AI to enhance emotional expression and feedback for school-aged children and adolescents (seo2024chacha; kim2018can; ibrahim2025designing). Meanwhile, other research focuses on providing specialized learning support for children with clinical needs, such as autism (tang2024emoeden; ghiotti2023prototyping; sharma2016promoting) or trauma (grassl2024coding). However, insufficient attention is given to children aged 3 to 6, highlighting a critical gap in HCI-driven emotional education.
Compared to older children, 3- to 6-year-olds experience emotions more immediately, intensely, and situationally while having limited verbal and self-regulation skills (sun2024exploring). These traits make parental guidance central for fostering emotional competence (gottman1996parental; hightower2019exploring). However, many parents face barriers such as time constraints and a lack of knowledge in addressing their children’s emotional needs (ahmadpour2023understanding; quan2025parents; BROWN2022126). Existing tools frequently overlook these developmental traits, assuming children can engage with AI independently and failing to empower parents in their crucial coaching role (sharma2025robot).
@@ 2.2. Parent-AI Collaborative Coaching System
AI-centered systems position AI as a “teaching tool” that interacts directly with children, offering guidance to help children (wang2024chaitok; wang2022don; karunaratna2024evaluation). For example, systems may generate personalized learning content (tang2024emoeden), support children in learning math language via story-based interaction (xu2023mathkingdom), co-create stories to boost vocabulary and feedback quality (zhang2024mathemyths), or generate context-based prompts to aid language expression (seo2024chacha). These systems emphasize direct AI-child interaction.
Parent-centered systems emphasize parents as primary educators, with AI as a supportive tool that does not interact directly with the child. These studies often use interviews or co-design methods to explore parents’ needs and challenges (dumaru2023after; cingel2017parents; nikkhah2021helping; hiniker2016not; mazmanian2017okay; wang2021protection). For example, research shows how parents co-use digital media with children (yu2024parent) and welcome daily prompts for reflection (ibrahim2025designing). However, few systems support real-world implementation, especially regarding sustained engagement and feedback mechanisms.
Parent-AI collaborative systems integrate the strengths of both parties through a division of labor. Parents can act as information providers, content configurators, facilitators, and evaluators (slovak2016scaffolding; chen2020new; yarosh2013almost). Meanwhile, AI is used to facilitate communication and storytelling (choi2024aacesstalk; zhang2022storybuddy), structure collaborative tasks and routines (jonas2023supporting; xiao2012supporting), and integrate smart devices into daily life (theofanopoulou2022exploring; chan2017wakey). These systems establish a three-way interaction model among the parent, child, and AI.
@@ 2.3. Modeling Children’s Emotional States
Personalized instruction is crucial in modern education and children’s emotional development (bernacki2021systematic; dewey2024democracy). Effective personalization requires modeling a child’s emotional state to enable targeted interventions. We examine the challenges of such modeling through two aspects: content design and assessment methods.
Regarding content design, a classic assessment for children’s emotional development is the Test of Emotion Comprehension (TEC) (pons2004emotion; albanese2006children). TEC identifies nine key elements for evaluating children’s emotional competence, which have also been widely acknowledged in prior research, including: emotion recognition (bullock1985further), emotional expression (barden1980children), emotion external cause understanding (harris1989young), emotion–belief link (fonagy1997relationship), emotion reminder effect (harris1983children), emotional masking (gardner1988japanese), mixed emotional understanding (arsenio1995children), emotional morality (harter1989developmental), and emotion regulation (altshuler1989developmental).
Regarding assessment methods, traditional approaches like parent-reported or teacher-reported questionnaires offer high ecological validity but suffer from subjectivity and inefficiency (denham1997parental; rothbart2004temperament). In contrast, self-report instruments like TEIQue-CF (mavroveli2008investigation) provide direct insight into children’s self-perceived emotional competence. However, for young children aged 3 to 6, limited cognitive and expressive abilities (wellman2004scaling) can lead to unreliable results. More recently, emerging AI systems using chatbots or games show promise for automated observation (santos2020therapist; nicolaidou2022gamified). However, these systems often struggle to interpret subtle and developmentally specific cues, especially when the data types available are limited due to privacy concerns (calvo2010affect).
@@ 3. STUDY PROCEDURE
Phase 2: Iterative design and implementation. To mitigate these challenges, we developed three core modules of PACEE: Parent-Child Adventure Game ($C_{1}$), Parental Feedback Report ($C_{2}$), and Child Emotional Modeling ($C_{3}$). PACEE follows Gottman’s Five Steps of Emotion Coaching framework, the PERMA model, and the design details advised from formative study interviews. Section 5 describes the design and implementation details.

### v2

@@ Abstract.
Emotion education is a crucial lesson for children aged 3 to 6. However, existing technologies primarily focus on promoting emotion education from the child’s perspective, often neglecting the central role of parents in guiding early childhood emotion development at home. In this work, we conducted co-design sessions with five experienced kindergarten teachers and five parents to identify parental challenges and the roles that AI can play in family emotion education. Guided by these insights, we developed PACEE, an assistant for supporting parent-AI collaborative emotion education. PACEE enables parents to engage in conversations about common emotional scenarios, with multiple forms of AI support to address parents’ challenges. It combines insights from parents and AI to model children’s emotional states and delivers personalized, parent-mediated guidance. In a user study involving 16 families, we found that PACEE significantly enhances parent-child engagement, encourages more in-depth emotional communication, and improves the parental experience. Our findings advance emotion coaching guidelines for family education in the era of generative AI, offering valuable insights for designing AI-supported, parent-centered family education systems.
@@
$c$s: Human-centered computing Collaborative interaction
@@ 1. INTRODUCTION
Emotion education plays a critical role in the developmental years of children aged 3 to 6. According to developmental psychology, during this period, children begin to understand various emotions, such as anger, fear, and sadness, and learn to express and regulate these feelings (Bartlett et al., 1982; Saarni, 1999; Wulandari, 2021).
Recent studies have explored AI-based systems that help children recognize, express, and manage their emotions. However, these systems primarily target children aged 8 and above, as they rely on independent conversations with LLM-based agents — for example, expressing emotions through conversations (Seo et al., 2024; Tang et al., 2024; Shen et al., 2025). Such designs are not well-suited for younger children aged 3 to 6, since at this stage, parents play a foundational role in shaping children’s emotional development by guiding their early emotional learning at home (Fadlillah and Fauziah, 2022; Alegre, 2011). Unlike older children, those aged 3 to 6 require parental support to engage with technology (Toran et al., 2024; Morawska et al., 2023) and rely on their parents for emotional security and co-regulation (Corwin, 2012; Brock and Kochanska, 2016).
However, providing effective emotion education can be challenging for parents, as most parents lack expertise in child psychology and strong communication skills (Sanders, 1999; Saarni, 1999; Cole et al., 1994). In practice, parents often encounter difficulties, such as misreading children’s ambiguous emotional signals (Castro et al., 2015) or failing to support emotional regulation through conversations (Meng et al., 2020).
@@
We developed PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education for children aged 3 to 6. PACEE helps parents engage in emotion education with their children through interactive adventure games. PACEE consists of three key modules, each addressing one of the challenges mentioned above: (1) Parent-Child Adventure Game. In each adventure game, parents and their children engage in a five-stage conversation that operationalizes Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997). The AI system provides parents with real-time guidance, such as actionable suggestions and progress tracking, thus solving $C_{1}$. (2) Parental Feedback Report. This module provides parents with feedback after each adventure game. The feedback includes a comprehensive review of past parent-child conversations and actionable suggestions for future interactions, effectively addressing $C_{2}$. (3) Child Emotional Modeling. This module creates an emotional profile for each child from two sources: the primary source being parents’ long-term understanding, elicited through an AI interviewer, supplemented by text-based analysis of parent-child conversation data. The emotional profile is then used to provide tailored guidance to parents during the adventure game. After each adventure game, it also offers a comparison between parents’ understanding of their child’s emotions and the AI’s analysis, thereby effectively addressing $C_{3}$.
To evaluate PACEE’s effectiveness in supporting family emotion education, we conducted a user study with 16 pairs of parents and children. Compared to a baseline (LLM-assisted) method, PACEE significantly fosters meaningful parent-child interactions and in-depth emotion education. Furthermore, with PACEE, children show greater engagement and richer emotional expression during conversations. Meanwhile, parents (1) improve their expertise and confidence in communication, (2) adopt more effective approaches in emotional education, and (3) gain a better understanding of their children’s emotional states.
@@
  • Through a formative study, we identified three key challenges in family emotion education, the role of AI in supporting this process, and principles for AI-assisted emotion education from the experts’ perspectives.
  • We present PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education. PACEE collaborates with parents to build children’s emotional profiles and provides personalized, parent-mediated guidance.
  • Through a user study with 16 families, we validated PACEE’s effectiveness in enhancing family emotion education. Our work advanced emotion coaching theory in both family and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
@@ 2.1. Emotion Coaching Techniques and Theories
Emotion coaching is a supportive parenting technique that helps children understand and regulate their emotions, and has been widely studied in the fields of parenting and child development (Fu et al., 2022; Havighurst and Kehoe, 2021). A key framework widely applied in HCI (Havighurst and Kehoe, 2021) is Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997; Gottman, 2005), which includes: (1) Be aware of the child’s emotion, (2) Recognize the child’s expression of emotion as a perfect moment for intimacy and teaching, (3) Listen with empathy and validate the child’s feelings, (4) Help the child learn to label their emotions with words, and (5) Set limits when helping the child to solve problems. This guideline is often complemented by principles from positive psychology (Baker et al., 2017; Fan et al., 2018), such as the PERMA model (Seligman, 2010). The PERMA model outlines five elements of well-being (Positive Emotion, Engagement, Relationships, Meaning, and Achievement) and has been applied in HCI areas such as well-being-centered design (Kovich et al., 2023), persuasive technologies (Nebrida and Dullas, 2018), and digital mental health (Chisale and Phiri, 2022).
Current HCI implementations of emotion coaching primarily target older children or specific clinical populations. Research has explored using AI to enhance emotional expression and feedback for school-aged children and adolescents (Seo et al., 2024; Kim et al., 2018; Ibrahim et al., 2025). Meanwhile, other research focuses on providing specialized learning support for children with clinical needs, such as autism (Tang et al., 2024; Ghiotti et al., 2023; Sharma et al., 2016) or trauma (Graßl and Fraser, 2024). However, insufficient attention is given to children aged 3 to 6, highlighting a critical gap in HCI-driven emotional education.
Compared to older children, 3- to 6-year-olds experience emotions more intensely and situationally, while having limited verbal and self-regulation skills (Sun et al., 2024). These traits make parental guidance central for fostering emotional competence (Gottman et al., 1996; Hightower et al., 2019). However, many parents face barriers such as time constraints and a lack of knowledge in addressing their children’s emotional needs (Ahmadpour et al., 2023; Quan et al., 2025; Brown et al., 2022). Existing tools frequently overlook these developmental traits, assuming children can engage with AI independently and failing to empower parents in their crucial coaching role (Sharma et al., 2025).
@@ 2.2. Parent-AI Collaborative Coaching System
AI-centered systems position AI as a “teaching tool” that interacts directly with children, offering guidance to help them (Wang et al., 2024, 2022; Karunaratna et al., 2024). For example, systems may generate personalized learning content (Tang et al., 2024), support children in learning math language via story-based interaction (Xu et al., 2023b), co-create stories to boost vocabulary and feedback quality (Zhang et al., 2024), or generate context-based prompts to aid language expression (Seo et al., 2024). These systems emphasize direct AI-child interaction.
Parent-centered systems emphasize parents as primary educators, with AI as a supportive tool that does not interact directly with the child. These studies often use interviews or co-design methods to explore parents’ needs and challenges (Dumaru and Al-Ameen, 2023; Cingel and Piper, 2017; Nikkhah et al., 2021; Hiniker et al., 2016; Mazmanian and Lanette, 2017; Wang et al., 2021). For example, research shows how parents co-use digital media with children (Yu et al., 2024) and welcome daily prompts for reflection (Ibrahim et al., 2025). However, few systems support real-world implementation, particularly in terms of sustained engagement and feedback mechanisms.
Parent-AI collaborative systems integrate the strengths of both parties through a division of labor, leveraging the complementary capabilities of each. Parents can act as information providers, content configurators, facilitators, and evaluators (Slovák et al., 2016; Chen et al., 2020; Yarosh et al., 2013). Meanwhile, AI is used to facilitate communication and storytelling (Choi et al., 2024; Zhang et al., 2022), structure collaborative tasks and routines (Jonas et al., 2023; Xiao and Martin, 2012), and integrate smart devices into daily life (Theofanopoulou and Slovak, 2022; Chan et al., 2017). These systems establish a three-way interaction model among the parent, child, and AI.
@@ 2.3. Modeling Children’s Emotional States
Personalized instruction is crucial in modern education and children’s emotional development (Bernacki et al., 2021; Dewey, 2024). Effective personalization requires modeling a child’s emotional state to enable targeted interventions. We examine the challenges of such modeling through two aspects: content design and assessment methods.
Regarding content design, a classic assessment for children’s emotional development is the Test of Emotion Comprehension (TEC) (Pons et al., 2004; Albanese et al., 2006). TEC identifies nine key elements for evaluating children’s emotional competence, which have also been widely acknowledged in prior research, including: emotion recognition (Bullock and Russell, 1985), emotional expression (Barden et al., 1980), emotion external cause understanding (Harris et al., 1989), emotion-belief link (Fonagy et al., 1997), emotion reminder effect (Harris, 1983), emotional masking (Gardner et al., 1988), mixed emotional understanding (Arsenio and Lover, 1995), emotional morality (Harter and Whitesell, 1989), and emotion regulation (Altshuler and Ruble, 1989).
Regarding assessment methods, traditional approaches, such as parent-reported or teacher-reported questionnaires, offer high ecological validity but suffer from subjectivity and inefficiency (Denham et al., 1997; Rothbart et al., 2004). In contrast, self-report instruments like TEIQue-CF (Mavroveli et al., 2008) provide direct insight into children’s self-perceived emotional competence. However, for young children aged 3 to 6, limited cognitive and expressive abilities (Wellman and Liu, 2004) can lead to unreliable results. More recently, emerging AI systems, such as those using chatbots or games, show promise for automated observation (Santos et al., 2020; Nicolaidou et al., 2022). However, these systems often struggle to interpret subtle and developmentally specific cues, especially when the data types available are limited due to privacy concerns (Calvo and D’Mello, 2010).
@@ 3. STUDY PROCEDURE
Phase 2: Iterative design and implementation. To mitigate these challenges, we developed three core modules of PACEE: Parent-Child Adventure Game ($C_{1}$), Parental Feedback Report ($C_{2}$), and Child Emotional Modeling ($C_{3}$). PACEE operationalizes Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997) and follows the design details advised from formative study interviews. Section 5 describes the design and implementation details.